OpenStudy (anonymous):

Write an equation in slope-intercept form of the line through point P(6, –1) with slope 4. @PeachRings

OpenStudy (anonymous):

13 more questions and I promise I will leave you alone

OpenStudy (anonymous):

13? hun i gotta go study for my nuclear chem exam

OpenStudy (anonymous):

ok i'm pretty sure someone else will help just let me know when you gotta go

OpenStudy (anonymous):

right meow

OpenStudy (anonymous):

ok bye thanks for help

OpenStudy (anonymous):

@NateRobinson

OpenStudy (anonymous):

@texaschic101 @Luigi0210 @OrangeMaster @david111 @dumbcow

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

Write an equation in slope-intercept form of the line through point P(6, –1) with slope 4. A. y = 4x – 1 B. y = 4x – 25 C. y + 1 = 4(x – 6) D. y + 6 = 4(x – 1)

OpenStudy (anonymous):

its B.

OpenStudy (anonymous):

yes that explains it very well @david111

OpenStudy (anonymous):

u welcome

OpenStudy (anonymous):

@NateRobinson thank you too

OpenStudy (anonymous):

AnyTime...